# Dynamical similarity analysis uniquely captures how computations develop in RNNs

### v2

@@ Abstract
Methods for analyzing representations in neural systems have become a popular tool in both neuroscience and mechanistic interpretability. Having measures to compare how similar activations of neurons are across conditions, architectures, and species, gives us a scalable way of learning how information is transformed within different neural networks. In contrast to this trend, recent investigations have revealed how some metrics can respond to spurious signals and hence give misleading results. To identify the most reliable metric and understand how measures could be improved, it is going to be important to identify specific test cases which can serve as benchmarks. Here we propose that the phenomena of compositional learning in recurrent neural networks (RNNs) would allow us to build a test case for dynamical representation alignment metrics. By implementing this case, we show it allows us to test whether metrics can identify representations which gradually develop throughout learning and probe whether representations identified by metrics are relevant to the actual computations executed within the network. By building both an attractor- and RNN-based test case, we can show that the recently proposed Dynamical Similarity Analysis (DSA) is more noise robust and identifies behaviorally relevant representations significantly more reliably than prior metrics (Procrustes, CKA). We also show how such test cases can be used beyond evaluating metrics to study new architectures directly. Specifically, we tested DSA in modern (Mamba) state space models, where results suggest that, in contrast to RNNs, these models may not exhibit or require changes in their recurrent dynamics due to their expressive hidden state. Overall, we develop test cases that can demonstrate how DSA’s increased ability to detect dynamical motifs gives it a superior ability to identify the ongoing computations in RNNs and elucidate how tasks are learned in networks.
@@ 1 Introduction
Both neuroscience and mechanistic interpretability want to understand how neural network solve the problems they face (He et al., 2024; Lindsay & Bau, 2023; Vilas et al., 2024). As the observed networks grow and the tasks to be understood become more complex, there is a need to have tools which allow to easily compare different architectures or parts of architectures across a wide array of task conditions and training parameters, at scale. One approach which has become popular is the use of representational alignment metrics which allow to conduct such comparative systems analyses on the activations observed across network (Sucholutsky et al., 2023) (Fig. 1). In mechanistic interpretability they can be used to compare how information is represented across layers (Raghu et al., 2021) and in neuroscience they allow to test under which conditions artificial neural networks work like the ventral stream of the brain (Kietzmann et al., 2019).
@@
While much of the literature is focused on comparison of static representations, meaning the activations observed in networks to non-dynamic stimuli like images, there has been an increasing interest in capturing the similarity of the dynamics of representations, so how representations change over the time course of a problem while the network is computing the correct response (Ju & Bassett, 2020) (Fig. 1). This is important due to the dynamic nature of both neuroscientific data (Vyas et al., 2020) and recurrently computing network models (Gu & Dao, 2023). Recent metrics like Dynamic Similarity Analysis (DSA) (Ostrow et al., 2024) address this challenge.
@@
In this work we propose that we can build on cases of well-understood simulated attractors and learning in recurrent neural networks (RNNs) to build test cases that assess how well metrics respond to dynamics of representations while also testing whether identified representations are connected to the computations and behaviors performed by the neural systems. Specifically, we construct two test cases: The first, based on simulated attractor dynamics, allows us to assess in a fully-controlled manner how robust different similarity measures are in identifying dynamically motifs commonly observed during computations of stateful neural networks, in the presence of noise and when faced with compositionally combined dynamical motifs. This fully controlled setup allows us to make normative statements about how a perfect ratio-like measure should numerically behave. The second test case is based on recent findings showing that RNNs learn to solve a compound task, made up of multiple subtask, by compositionally combining the representational dynamics of all subtasks (Driscoll et al., 2024). This test case allows us to make ordinal predictions for how similar representations should be across the compositional learning process and at the same time creates the opportunity to directly link gradually developing representations to gradually developing computational abilities.
With these test cases we can compare three metrics used to compare dynamic representations, namely Procrustes transformation (Procrustes), Centered Kernel Alignment (CKA), and Dynamic Similarity Analysis (DSA). This leads to our main contributions:
  • DSA shows better noise-robustness and a more ratio-like response to our attractor test case
  • DSA is the only measure which responds correctly to the compositional learning of the RNN-based test case
  • DSA is the only measure that can link representations to computations by showing that increasingly similar representations are predictive of increasingly similar behavior in the RNN-based test case
  • When applying DSA and our test case to state space models (SSMs) the results suggest that Mamba (Gu & Dao, 2023) might not require changes to their recurrent dynamics during learning of simple tasks, due to their expressive hidden state
@@ 2 Related work
A lot of progress has been made on using representational alignment metrics to both compare and induce similarities across models and empirical data (Sucholutsky et al., 2023; Williams et al., 2021). With representational metrics gaining popularity, researchers have identified the need to properly compare metrics to understand their shortcoming (Dujmović et al., 2023; Soni et al., 2024). Recent work on proposing ways to benchmark static representational alignment metrics tries to give us the ability to make an informed decision which metric is best to use (Klabunde et al., 2024; Ahlert et al., 2024). The goal is not necessarily to find “the ultimate best metric” but instead to highlight in which context to rely on a specific metric or sets of metrics, similar specialized statistical tests.
Our work specifically focuses on comparing dynamic representations (as opposed to static representations; see Fig. 1). While researchers have been interested in using representational alignment metrics for dynamic use cases, they usually relied on extending static metrics to also be applied to dynamic use cases (Lin et al., 2019; Williams et al., 2018; Kietzmann et al., 2019; Maheswaranathan et al., 2019; Cloos et al., 2024). With increasing interest in dynamic metrics, we have now also seen new approaches being developed to compare dynamical representations (Redman et al., 2024; Ostrow et al., 2024; Chen et al., 2024). Generally speaking, being specialized for dynamical representations means that metrics can actively capture the momentum of traces in instead of just their shapes. For example, DSA (Ostrow et al., 2024) compares the temporal dynamics of two systems by embedding them in a high-dimensional space and using a modified Procrustes analysis to assess the similarity of their vector fields. Another example is Diffeomorphic Vector Field Alignment (Chen et al., 2024) which evaluates the similarity by learning a nonlinear coordinate transformation that aligns vector fields, providing a measure of functional similarity through orbital equivalence. While new metrics are being developed and used (Eisen et al., 2024), the question of how to decide when to use which specific metric is still an open problem for the case of dynamic metrics.
@@
In the following we address these two trends in the literature: we extend the idea of benchmarking representational alignment metrics to the dynamic case, and we do so in a way that allows us to test if the representations connect to the system’s computations by testing whether gradually evolving representations link to gradually developing computational abilities.
@@ 3 Methods
3.1 Metrics
For the following analyses we use the three metrics Centered Kernel Alignment (‘CKA’, Kornblith et al. (2019)), Procrustes Transformation (‘Procrustes’, Cloos et al. (2024)) and Dynamical Similarity Analysis (‘DSA’, Ostrow et al. (2024)). CKA and Procrustes both consider the dynamics as static traces and try to either align the two traces through optimal geometric transformation (Procrustes) or measure the similarity between the kernel matrices of the feature spaces (CKA). In contrast, DSA considers the traces as dynamic and projects them into high dimensional linear spaces where it compares the transition matrix of traces (Fig. 1). We use DSA with the hyperparameter number of delays = 33 and delay interval = 6. Appendix 6.1 explains how parameters were set.
3.2 RNN / SSM training and analysis
For every condition we trained 72 RNNs, one for each combination of the following hyperparameters: architecture (Leaky RNN, Leaky GRU), activation function (relu, softplus, tanh), hidden size (128, 256), learning rate (1e-2, 1e-3), batch size (64, 128, 256). Hyperparameters were chosen based on Driscoll et al. (2024). One epoch of training data consisted of 10000 trials for each task. When a network was pretrained on multiple tasks, we trained the tasks sequentially within each epoch (e.g. for each epoch, first train on task A, then B). Networks were trained in a supervised fashion (with Adam optimizer) until they reached 99% accuracy on each task or for 50 epochs, depending on what was reached earlier. We only analyzed the networks that managed to fully learn the master task to 99% accuracy, which corresponded to 91% of all models.
The training of the Mamba SSMs (Gu & Dao, 2023) works in the same way as RNNs. For every condition we trained 64 models, one for each combination of the following hyperparameters: number of layers (1, 2), hidden dimensions (8, 16), learning rate (1e-2, 5e-3, 1e-3, 5e-4), batch size (16, 32, 64, 128). We only analyzed the networks that managed to fully learn the master task, which corresponded to 99% of all models. We used the implementation by Torres-Leguet (2024).
3.3 RNN tasks
For the RNN-based test case, we use task implementations based on Neurogym (Molano-Mazon et al., 2022). Network inputs are the standard task inputs alongside an additional task identifier which identifies the four tasks used (A, B, C, M) through one-hot encoding. We use three subtasks. In task A and B (Pro and Anti Task) the model receives two continuous numbers as separate inputs, with time-varying Gaussian noise. It needs to decide which input is higher (Pro) or lower (Anti) stimulus. Task C (Delay Task) is the same as A but with an additional delay. These are combined to form a compound ‘Master’ task (M, DelayAnti Task), where networks need to determine which stimulus was lower, after a delay period. Each task consisted of a stimulus presentation period (200 time steps) and a choice period (25 time steps), and an optional delay period as described above. The duration of the delay is variable during training (25, 50 or 75 time steps) but fixed during testing (100 time steps). We only analyze the hidden states during the stimulus presentation period, keeping the first twenty principal component after centering and normalization. While networks are optimized for correct fixation during stimulus presentation and delay periods, as well as correct choices after the fixation, when we report accuracy, we report it based on the choice made by the network during the response period at the end of the trial, weighted towards the last time step.
@@ 4 Experiments
Through the following experiments we want to test whether representational alignment metrics can capture the representations and computations of stateful artificial neural networks. We specifically compare the performance of CKA, Procrustes, and DSA. The first test case is focused on attractor dynamics usually observed to be underlying computations, followed by analyses directly linking identified dynamics with computations of trained RNNs. We close by applying metrics to analyze newly developed SSMs. The code used in this study is available at GitHub link for reproducibility and further exploration.
4.1 Testing for ‘ratio-response’ of metrics to noisy combined attractors
In this first section, we want to test how well metrics capture attractor dynamics in the presence of noise and when multiple attractor dynamics are combined compositionally. For this we simulate attractor dynamics using Lorenz attractors ($\sigma=10$; $\beta=2.667$). This first set of simulations is close to what is usually examined when validating new dynamic representation measures but here we focus on the specific abilities which underlie the complex test case introduced later. This allows us to test for these abilities in a simpler setting where we can also make normative statements about how metrics need to behave to act as a ratio. A measure behaves as a ratio if it preserves order, has equal intervals between values, and has a meaningful zero point, allowing for meaningful ratios between any two measurements.
Figure 2: DSA shows better noise robustness and interval scaling. (a) Outline of model specifications to test both noise robustness (Model 1 vs. Model 2) and ratio-scaling (Model 2 vs. Model 3). (b) Results of noise robustness comparisons. (c) Results of ratio scaling comparisons. All shadings are standard errors.
@@ 4.1 Testing for ‘ratio-response’ of metrics to noisy combined attractors
All metrics can identify basic attractor motifs (see Appendix 6.2) but we how well metrics can identify attractors in the presence of noise? For this, we construct two ‘noise + attractor’ combinations (‘models’) (Fig. 2a). Model 1 combines one attractor sample (Attractor A) with one sample of Gaussian noise (Noise A). Model 2 combines the same attractor sample (Attractor A) with a different sample of Gaussian noise (Noise 2). We now gradually reduce the noise level across both noise samples in three steps (‘Epochs’), by 50% each time, and compare Model 1 with Model 2 within each Epoch. We sample 7 sections of different Lorenz attractors (see Appendix 6.2), and each is used once as ‘Attractor A’. Each attractor is used 200 times with different noise samples. Hence we run a total of 1400 comparisons. Results are depicted in Fig. 2b. As we want metrics to respond to the true underlying dynamic and ignore the noise, we want to see that across epochs, the similarity is a horizontal line, so not changing in response to changing noise levels. This is measured by a low dissimilarity gap (difference in dissimilarity between epoch 1 and 3). DSA shows the lowest response to noise ($1.7\times 10^{-2}$ vs $4.5\times 10^{-2}$ [CKA] vs $3.9\times 10^{-2}$ [Procrustes]).
@@
We also want to analyze whether metrics respond ratio-like to combined attractors. For this we assemble ‘Model 3’ which combines Attractor A with another Attractor B and a new sample of Gaussian noise (Noise 3). For Model 3, we gradually decrease both the noise and the relative relevance of Attractor B. We use the same 7 attractor samples as above, so we get 42 attractor pairings where Attractor A and Attractor B are not the same. Each attractor pairing is used 200 times with different noise samples. Hence we run a total of 8400 comparisons. We would like to see a linear decrease in dissimilarity across epochs as measured by the linearity measure. We assess the linear decrease of dissimilarity by computing the R² of the linear fit of a regression over the dissimilarity predicted by the epoch (Fig. 2c). DSA shows the most linear decrease ($0.99$ vs $0.96$ [CKA] vs. $0.97$ [Procrustes]). Note that for real ratio-like behaviors, metrics should decrease linearly and conclude on the value of 0 for full similarity. Fig. 2b highlights that while DSA shows the least change to noise, it is furthest away from true zero. This results in DSA having a lower dissimilarity gap in Fig. 2c. Both Procrustes and CKA get closer to true zero. This highlights that DSA might need additional normalization to function like a proper ratio.
4.2 Testing ordinal response of metrics to compositional learning setup

### v3

# Dynamical similarity analysis can identify compositional dynamics developing in RNNs

@@ Abstract
Methods for analyzing representations in neural systems have become a popular tool in both neuroscience and mechanistic interpretability. Having measures to compare how similar activations of neurons are across conditions, architectures, and species, gives us a scalable way of learning how information is transformed within different neural networks. In contrast to this trend, recent investigations have revealed how some metrics can respond to spurious signals and hence give misleading results. To identify the most reliable metric and understand how measures could be improved, it is going to be important to identify specific test cases which can serve as benchmarks. Here we propose that the phenomena of compositional learning in recurrent neural networks (RNNs) allows us to build a test case for dynamical representation alignment metrics. By implementing this case, we show it enables us to test whether metrics can identify representations which gradually develop throughout learning and probe whether representations identified by metrics are relevant to computations executed by networks. By building both an attractor- and RNN-based test case, we show that the new Dynamical Similarity Analysis (DSA) is more noise robust and identifies behaviorally relevant representations more reliably than prior metrics (Procrustes, CKA). We also show how test cases can be used beyond evaluating metrics to study new architectures. Specifically, results from applying DSA to modern (Mamba) state space models, suggest that, in contrast to RNNs, these models may not exhibit changes to their recurrent dynamics due to their expressiveness. Overall, by developing test cases, we show DSA’s exceptional ability to detect compositional dynamical motifs, thereby enhancing our understanding of how computations unfold in RNNs.
@@ 1 Introduction
Both neuroscience and mechanistic interpretability aim to understand how neural networks solve the problems they face (He et al., 2024; Lindsay & Bau, 2023; Vilas et al., 2024). As the architectural complexity of such models increases, there is a growing need to have tools which allow us to compare models across a wide array of task conditions and training parameters, at scale. One approach which has become popular is the use of representational alignment metrics which allow to conduct comparative analyses on the activations observed across networks (Sucholutsky et al., 2023) (Fig. 1). In mechanistic interpretability they can be used to compare how information is represented across layers (Raghu et al., 2021) and in neuroscience they allow us to test under which conditions neural networks work like specific brain regions (e.g. ventral stream of the brain; Kietzmann et al. (2019)).
@@
While much of the literature is focused on static representations, meaning the activations observed in networks to non-dynamic stimuli like images, there has been an increasing interest in capturing the similarity of the dynamics of representations, so how representations change over the time course of a problem while the network is computing the correct response (Ju & Bassett, 2020) (Fig. 1). This is important due to the dynamic nature of both neuroscientific data (Vyas et al., 2020) and recurrently computing network models (Gu & Dao, 2023). Recent metrics like Dynamic Similarity Analysis (DSA) (Ostrow et al., 2024) address this challenge.
@@
In this work, we propose that we can use well-understood cases of simulated attractors and learning in recurrent neural networks (RNNs) to build test cases that assess how well metrics respond to dynamical representations while also testing whether such representations can be connected to the computations performed by networks. Specifically, we construct two test cases: The first, based on simulated attractor dynamics, allows us to assess in a controlled manner how robust similarity measures are in identifying dynamical motifs commonly observed during computations of stateful neural networks, in the presence of noise and when dynamical motifs are compositionally combined. This controlled setup allows us to evaluate metrics based on quantitative predictions. The second test case, builds on findings showing that RNNs learn to solve a compound task, made up of multiple subtasks, by compositionally combining the representational dynamics of subtasks (Driscoll et al., 2024). This test case allows us to make ordinal predictions for how similar representations should become across compositional learning and creates the opportunity to link gradually developing representations with the resulting computations.
With these test cases, we can compare three metrics used to compare dynamic representations, namely Procrustes transformation (Procrustes), Centered Kernel Alignment (CKA), and Dynamic Similarity Analysis (DSA). Our main contributions are showing that:
  • DSA shows better noise-robustness and ability to identify combined dynamics
  • Out of the metrics we tested, DSA is the only measure which responds correctly to the compositional learning of the RNN-based test case
  • DSA is the only measure that can link representations to computations by showing that increasingly similar representations are predictive of increasingly similar behavior in RNNs
  • When applying DSA to state space models (SSMs) our results suggest that Mamba (Gu & Dao, 2023) might not require changes to dynamics during learning of simple tasks
@@ 2 Related work
A lot of progress has been made on using representational alignment metrics to both compare and induce similarities across models and empirical data (Sucholutsky et al., 2023; Williams et al., 2021). With representational metrics gaining popularity, researchers have identified the need to properly compare metrics to understand their shortcomings (Dujmović et al., 2023; Soni et al., 2024). Efforts to propose methods for benchmarking static representational alignment metrics aim to enable informed decisions about which metric is most suitable for a given application (Klabunde et al., 2024; Ahlert et al., 2024). Therefore, the goal is not necessarily to find “the best metric” but instead to highlight in which context to rely on a specific metric, similar to specialized statistical tests.
Our work specifically focuses on comparing dynamic representations (as opposed to static representations; see Fig. 1). While researchers have been interested in using representational alignment metrics for dynamic use cases (Lin et al., 2019; Williams et al., 2018; Kietzmann et al., 2019; Maheswaranathan et al., 2019; Cloos et al., 2024), they usually relied on extending static metrics to also be applied to dynamic use cases. With growing interest in dynamic metrics, we are now seeing the development of new approaches to compare dynamical representations (Redman et al., 2024; Ostrow et al., 2024; Chen et al., 2024). Generally speaking, being specialized for dynamical representations means that metrics can actively capture the momentum of traces (i.e. the speed by which a system moves from point A in state space to point B in state space) in instead of just their shapes (i.e. only capturing whether the system moves from A to B at all). For example, DSA (Ostrow et al., 2024) compares the temporal dynamics of two systems by embedding them in a high-dimensional space and using a modified Procrustes analysis to assess the similarity of their vector fields. Another example is Diffeomorphic Vector Field Alignment (Chen et al., 2024) which evaluates the similarity by learning a nonlinear coordinate transformation that aligns vector fields, providing a measure of functional similarity through orbital equivalence. Although new metrics are being developed and applied (Eisen et al., 2024; Lipshutz et al., 2024), the challenge of determining when to use a specific metric remains an open question, especially in the context of dynamic metrics.
@@
In the following we address these two trends in the literature: we extend the idea of benchmarking representational alignment metrics to the dynamic case. By conceiving two test cases we can investigate whether representational metrics respond to the representations of a neural network that link to the network’s computations. We do this by testing whether gradually evolving representations link to developing computational abilities throughout learning. These two isolated test cases could form part of a larger benchmark for dynamic representation alignment metrics in the future.
@@ 3 Methods
Metrics
We use the three metrics: Centered Kernel Alignment (‘CKA’, Kornblith et al. (2019)), Procrustes Transformation (‘Procrustes’, Cloos et al. (2024)) and Dynamical Similarity Analysis (‘DSA’, Ostrow et al. (2024)). CKA and Procrustes were initially conceived for static representations and hence consider the dynamics as static traces and try to either align the two traces through optimal geometric transformation (Procrustes) or measure the similarity between the kernel matrices of the feature spaces (CKA). These metrics have shown promise in being adaptable to dynamic representations Cloos et al. (2024). In contrast, DSA is a metric specifically developed for dynamic representations and considers neuronal traces as dynamic, projecting them into high dimensional linear spaces where it compares the transition matrix of traces (Fig. 1). We use DSA with the hyperparameter number of delays = 33 and delay interval = 6. Appendix 6.1 explains how parameters were set.
RNN / SSM training and analysis
For every condition, we trained 72 Recurrent Neural Networks (RNNs) and 64 State Space Models (SSMs) using the Mamba architecture (Gu & Dao, 2023) and the implementation by Torres-Leguet (2024). The hyperparameters for the RNNs and SSMs were chosen based on Driscoll et al. (2024). The training setup is expressed in more detail in Appendix 6.1.
Tasks for RNN-based test case
We used Neurogym task implementations to train RNNs (Molano-Mazon et al., 2022). Different task versions required networks to compare two stimuli, fixate during presentation and delay periods, and choose the higher or lower stimulus based on task-specific rules. Section 6.3 explains how task versions are created. Appendix table 3 contains additional details. Network inputs included the standard task inputs alongside a one-hot task identifier for the current task (A, B, C, M).
@@ 4 Experiments
Through the following experiments we want to test whether representational alignment metrics can capture the compositional representations and computations of stateful artificial neural networks. We specifically compare the performance of CKA, Procrustes, and DSA. The first test case is focused on simulated attractor dynamics that mirror dynamics usually observed to be underlying computations in RNNs. In the second case we analyze how identified dynamics with computations of trained RNNs. We close by applying metrics to analyze newly developed SSMs. The code used in this study is available at GitHub link for reproducibility and further exploration. Appendix 6.1 provides additional information on the methods used for the metrics and the training schedules.
4.1 Metrics’ ability to identify compositionally-combined noisy attractors dynamics
First, we want to test how well metrics capture attractor dynamics in the presence of noise and when multiple attractor dynamics are combined compositionally. We see these as core skills needed for metrics to perform in the more complex RNN-base test case introduced later (6.3). To test for these abilities, we simulate attractor dynamics using Lorenz attractors ($\sigma=10$; $\beta=2.667$).
@@ 4.1 Metrics’ ability to identify compositionally-combined noisy attractors dynamics
All metrics can identify basic attractor motifs (see Appendix 6.2) but how well can metrics identify attractors in the presence of noise? For this, we construct two ‘noise + attractor’ combinations (‘models’) (Fig. 2a and Fig 8) where ”+” refers to a numerical addition and scaling back to the unit volume. Model 1 combines one attractor sample (Attractor A) with one sample of Gaussian noise (Noise A). Model 2 combines the same attractor sample (Attractor A) with a different sample of Gaussian noise (Noise 2). We now gradually reduce the noise level across both noise samples in three steps (‘Combinations’), by 50% each time, and compare Model 1 with Model 2 within each set of Combinations. We sample 7 sections of different Lorenz attractors (see Appendix 6.2), and each is used once as ‘Attractor A’. Each attractor is used 200 times with different noise samples. Hence we run a total of 1400 comparisons. Results are depicted in Fig. 2b. As we want metrics to respond to the true underlying dynamic and ignore the noise, we want to see that across epochs, the similarity is a horizontal line, so not changing in response to changing noise levels. This is measured by a low dissimilarity gap (difference in dissimilarity between epoch 1 and 3). DSA shows the lowest response to noise ($1.7\times 10^{-2}$ vs $4.5\times 10^{-2}$ [CKA] vs $3.9\times 10^{-2}$ [Procrustes]).
@@
To identify compositional and computationally-relevant dynamics, metrics not only need to be noise robust but specifically noise robust in the context of compositionally-combined attractor motifs. To test for this ability, we assemble ‘Model 3’ which combines Attractor A with another Attractor B and a new sample of Gaussian noise (Noise 3). For Model 3, we gradually decrease both the noise and the relative amplitude of Attractor B. We use the same 7 attractor samples as above, so we get 42 attractor pairings where Attractor A and Attractor B are not the same. Each attractor pairing is used 200 times with different noise samples. Hence we run a total of 8400 comparisons. How do we know that a metric responds correctly to the combination of attractors? Given that we use random combinations of attractors and noise, with a linearly decreasing influence of the additional attractor and noise, we would like to see a linear decrease in dissimilarity across epochs as measured by the linearity measure. We assess the linear decrease of dissimilarity by computing the R² of the linear fit of a regression over the dissimilarity predicted by the epoch (Fig. 2c). DSA shows the most linear decrease ($0.99$ vs $0.96$ [CKA] vs. $0.97$ [Procrustes]). Note that one might want metrics to behave like a ratio to a test case like this. A measure behaves as a ratio if it has equal intervals between values, and has a meaningful zero point, allowing for meaningful ratios between any measurements. In our case that would mean metrics should decrease linearly and show a value of 0 for full similarity. Fig. 2b highlights that while DSA shows the least change to noise, it is furthest away from true zero. This results in DSA having a lower dissimilarity gap in Fig. 2c. Both Procrustes and CKA get closer to true zero. This highlights that DSA might need additional normalization to function like a ratio.
These analyses based on simple and well-controlled attractor dynamics reveal that DSA might have a very slight advantage at identifying dynamical motifs in the presence of noise and when they are compositionally combined. From these analyses alone it is not possible to infer whether these very minor quantitative difference result in qualitatively different behavior of metrics in more complex environments. The next section will hence explicitly focus on a more complex test case.
Figure 2: DSA shows better noise robustness and identification of compositionally-combined dynamics. (a) Outline of model specifications to test both noise robustness (Model 1 vs. Model 2) and identification of combined dynamics (Model 2 vs. Model 3). (b) Results of noise robustness (c) compositional dynamics comparisons. All shadings are standard errors.
4.2 Metrics’ ability to identify compositional dynamics during over training in RNNs